# Matter- and magnetically-driven flavor conversion of neutrinos in magnetorotational collapses

### 09 Mar 2026 v1

@@ 1 Introduction
Neutrinos are also expected to have a non-zero magnetic dipole moment [42]. Lower bounds coming from theoretical arguments suggest $\mu\gtrsim\mathcal{O}(10^{-23}$$10^{-22})\mu_{B}$, with $\mu_{B}=e/(2m_{e})$ being the Bohr magneton. The most stringent bounds from terrestrial experiments are compatible with $\lesssim\mathcal{O}(10^{-11})\mu_{B}$, while astrophysical bounds give $\lesssim\mathcal{O}(10^{-12})\mu_{B}$, see Refs. [42, 43] and references therein. Hence, in the presence of large magnetic fields, neutrinos may experience magnetically-induced flavor conversion [44, 45, 46, 47, 48]. If neutrinos are Majorana particles, neutrino-antineutrino flavor conversion can take place [49, 50, 51, 42]. However, the physics of neutrino flavor conversion in magnetorotational collapses has been poorly explored.
@@ 2 Benchmark model
Figure 1: Radial profiles of the baryon density ($\rho$, top row), electron fraction ($Y_{e}$, middle row), and perpendicular magnetic-field strength ($B_{\perp}$, bottom row) at two post-bounce times ($t=0.3\,\mathrm{s}$ and $t=1\,\mathrm{s}$), shown along the pole ($\theta=0$ and $\varphi=0$, left column) and the equator ($\theta=\pi/2$ and $\varphi=0$, right column). The dotted lines in the bottom panel represent the output of the neutrino-magnetohydrodynamic simulation, but are unphysical and due to the progenitor model employed to initialize the simulation. The solid lines represent our interpolated profiles that smoothly bridge these unphysical gaps (see main text for more details).
@@ 2.1 Magnetorotational core collapse model
The electron fraction $Y_{e}$ is plotted in the middle row of Fig. 1. In the dense proto-neutron star interior, where electron degeneracy is high and neutrinos are trapped, $Y_{e}\simeq 0.3$. Moving outwards, a pronounced dip in $Y_{e}$ develops corresponding to the deleptonization layer, produced by efficient electron capture ($e^{-}+p\rightarrow n+\nu_{e}$), which dominates neutrino absorption and drives the local electron fraction to a minimum. Beyond this layer, absorption of electron neutrinos off neutrons ($\nu_{e}+n\rightarrow p+e^{-}$) outweighs electron capture, pushing the composition toward a more symmetric proton-to-neutron ratio. As a result, $Y_{e}$ grows sharply, exceeding $0.5$ in some emission directions and for some post-bounce times. This trend is visible along the polar direction at $t=1$ s; proton-rich regions also develop along other emission directions.
The bottom row of Fig. 1 represents $B_{\perp}$, the component of the magnetic field relevant for the neutrino–antineutrino conversion discussed in Sec. 3.2. The inner regions exhibit strong magnetic-field strength ($B_{\perp}\simeq 10^{15}$$10^{16}$ G in the proximity of the proto-magnetar), reflecting the combined effect of flux compression and rotational winding. At larger radii, the field gradually decreases with radius and displays significant directional variability. At intermediate and large radii ($r\simeq 10^{4}$ km and just below $10^{5}$ km), the dotted curves show the simulation output exhibiting localized gaps, where the magnetic field artificially drops to zero [58]. These features are not physical; they arise from the progenitor model (used to initialize our benchmark simulation), where the magnetic field profile is prescribed in radiative regions only, while convective zones have $B=0$. As a consequence, the magnetorotational simulation retains regions of vanishing field in the vicinity of the convective shells. Since such gaps are not expected to occur and would artificially affect our analysis, we construct interpolated profiles (solid curves) that smoothly bridge these unphysical gaps. The interpolated profiles are the ones adopted in the rest of this work.
@@ 3.2 Neutrino-antineutrino conversion
Flavor conversion is enhanced when a pair (or more) of effective eigenstates becomes close to degenerate. Due to the different sign in the matter potential between neutrinos and antineutrinos, these flavor-changing chirality-flipping conversions can be resonant [49, 78]. This phenomenon is referred to as resonant spin-flavor precession [78, 49, 50]. Hereafter, we dub these transitions magnetically-driven resonant conversion (B-res) to avoid confusion, since MSW flavor conversion can be interpreted as resonant precession of a spin vector [81]. For simplicity, we assume equal magnetic moments for all flavors, i.e., $\mu_{e\mu}=\mu_{e\tau}=\mu_{\mu\tau}=\mu$. The ratio of magnetic moments $\mu_{e\mu}:\mu_{e\tau}:\mu_{\mu\tau}$ depends on the neutrino mass mechanism and underlying symmetries, see e.g. Refs. [82, 83, 84]. Accounting for differences among the magnetic moments (albeit plagued by large uncertainties to date) would only affect the adiabaticity of the resonances, but not their location.
@@ 3.3 Discussion
We focus on exploring the flavor conversion phenomenology for Majorana particles. However, for Dirac neutrinos, the chirality flip resulting from dipolar magnetic interactions implies left-handed neutrinos turn into right-handed ones (and right-handed antineutrinos into left-handed ones), which then evade detection via weak interactions. For Dirac neutrinos, matter can also enhance these conversions resonantly. The resonance location would be different with respect to the one investigated for Majorana neutrinos. As a result, neutrino magnetic dipole interactions may be responsible for a reduction of the detected flux for Dirac particles. This scenario may be degenerate with other uncertainties plaguing the expected flux at Earth, such as the core-collapse physics, the role of neutrino-neutrino interaction, and potential beyond-the-Standard-Model scenarios. For more details, we refer the reader to Ref. [42], where the phenomenology of B-res transitions is thoroughly addressed.
@@ 4 Flavor composition at Earth
The (anti)neutrino fluxes expected at Earth in NO, assuming adiabatic MSW transitions (Sec. 3.1) and allowing for generic values of the B-res jump probabilities $p_{i}$, are:

### 14 May 2026 v2

@@ 1 Introduction
Neutrinos are also expected to have a non-zero magnetic dipole moment [34]. Lower bounds coming from theoretical arguments suggest $\mu\gtrsim\mathcal{O}(10^{-23}$$10^{-22})\mu_{B}$, with $\mu_{B}=e/(2m_{e})$ being the Bohr magneton. The most stringent bounds from terrestrial experiments are compatible with $\lesssim\mathcal{O}(10^{-11})\mu_{B}$, while astrophysical bounds give $\lesssim\mathcal{O}(10^{-12})\mu_{B}$, see Refs. [34, 64] and references therein. Hence, in the presence of large magnetic fields, neutrinos may experience magnetically-induced flavor conversion [32, 21, 73, 85, 68]. If neutrinos are Majorana particles, neutrino-antineutrino flavor conversion can take place [7, 5, 1, 27, 34]. However, the physics of neutrino flavor conversion in magnetorotational collapses has been poorly explored.
@@ 2 Benchmark model
Figure 1: Radial profiles of the baryon density ($\rho$, top row), electron fraction ($Y_{e}$, middle row), and transverse magnetic-field strength ($B_{\perp}=\sqrt{B_{\theta}^{2}+B_{\varphi}^{2}}$, bottom row) at two post-bounce times ($t=0.3\,\mathrm{s}$ and $t=1\,\mathrm{s}$), shown along the pole ($\theta=0$ and $\varphi=0$, left column) and the equator ($\theta=\pi/2$ and $\varphi=0$, right column). The dotted lines in the bottom panel represent the output of the neutrino-magnetohydrodynamic simulation, but are unphysical and due to the progenitor model employed to initialize the simulation. The solid lines represent our interpolated profiles that smoothly bridge these unphysical gaps (see main text for more details).
@@ 2.1 Magnetorotational core collapse model
The electron fraction $Y_{e}$ is plotted in the middle row of Fig. 1. In the dense proto-neutron star interior, where electron degeneracy is high and neutrinos are trapped, $Y_{e}\simeq 0.3$. Moving outwards, a pronounced dip in $Y_{e}$ develops corresponding to the deleptonization layer, produced by efficient electron capture ($e^{-}+p\rightarrow n+\nu_{e}$), which dominates neutrino emission and drives the local electron fraction to a minimum. Beyond this layer, absorption of electron neutrinos off neutrons ($\nu_{e}+n\rightarrow p+e^{-}$) outweighs electron capture, pushing the composition toward a more symmetric proton-to-neutron ratio. As a result, $Y_{e}$ grows sharply at $t=0.3$ s and $t=1$ s, along both emission directions. As visible in Fig. 1, $Y_{e}$ exceeds $0.5$ only along the polar direction at $t=1$ s, indicating the development of proton-rich regions.
The bottom row of Fig. 1 represents $B_{\perp}$, the component of the magnetic field relevant for the neutrino–antineutrino conversion discussed in Sec. 3.2. The inner regions exhibit strong magnetic-field strength ($B_{\perp}\simeq 10^{15}$$10^{16}$ G in the proximity of the proto-magnetar), reflecting the combined effect of flux compression and rotational winding. At larger radii, the field gradually decreases with radius and displays significant directional variability, as visible from the differences between the polar and equatorial profiles.. At intermediate and large radii ($r\simeq 10^{4}$ km and just below $10^{5}$ km), the dotted curves show the simulation output exhibiting localized gaps, where the magnetic field artificially drops to zero [35]. These features are not physical; they arise from the progenitor model (used to initialize our benchmark simulation), where the magnetic field profile is prescribed in radiative regions only, while convective zones have $B=0$. As a consequence, the magnetorotational simulation retains regions of vanishing field in the vicinity of the convective shells. Since such gaps are not expected to occur and would artificially affect our analysis, we construct interpolated profiles (solid curves) that smoothly bridge these unphysical gaps. The interpolated profiles are the ones adopted in the rest of this work.
@@ 3.2 Neutrino-antineutrino conversion
Flavor conversion is enhanced when a pair (or more) of effective eigenstates becomes close to degenerate. Due to the different sign in the matter potential between neutrinos and antineutrinos, these flavor-changing chirality-flipping conversions can be resonant [7, 50]. This phenomenon is referred to as resonant spin-flavor precession [50, 7, 5]. Hereafter, we dub these transitions magnetically-driven resonant conversion (B-res) to avoid confusion, since MSW flavor conversion can be interpreted as resonant precession of a spin vector [45]. For simplicity, we assume equal magnetic moments for all flavors, i.e., $\mu_{e\mu}=\mu_{e\tau}=\mu_{\mu\tau}=\mu$. Here, $\mu_{\alpha\beta}$ denote the transition magnetic moments between neutrino flavors $\alpha$ and $\beta$. The ratio of magnetic moments $\mu_{e\mu}:\mu_{e\tau}:\mu_{\mu\tau}$ depends on the neutrino mass mechanism and underlying symmetries, see e.g. Refs. [86, 16, 30]. Accounting for differences among the magnetic moments (albeit plagued by large uncertainties to date) would only affect the adiabaticity of the resonances, but not their location.
@@ 3.3 Discussion
We focus on exploring the flavor conversion phenomenology for Majorana particles. However, for Dirac neutrinos, the chirality flip resulting from dipolar magnetic interactions implies left-handed neutrinos turn into right-handed ones (and right-handed antineutrinos into left-handed ones), which then evade detection via weak interactions. For Dirac neutrinos, matter can also enhance these conversions resonantly. The resonance location would be different with respect to the one investigated for Majorana neutrinos. As a result, neutrino magnetic dipole interactions may be responsible for a reduction of the detected flux for Dirac particles. This scenario may be degenerate with other uncertainties plaguing the expected flux at Earth, such as the core-collapse physics, the role of neutrino-neutrino interaction, and potential beyond-the-Standard-Model scenarios. For more details, we refer the reader to Ref. [34], where the phenomenology of B-res transitions is thoroughly addressed. Related resonant flavor conversion mechanisms for Dirac neutrinos induced by twisting magnetic fields have also been explored in Ref. [39].
@@ 4 Flavor composition at Earth
The (anti)neutrino fluxes expected at Earth in NO, assuming adiabatic MSW transitions (Sec. 3.1), accounting for the loss of coherence when travelling to Earth, and allowing for generic values of the B-res jump probabilities $p_{i}$, are: